OpenStudy (anonymous):

3x-9y=-12 6x+20y=-62

OpenStudy (anonymous):

3x-9y=-12 6x+20y=-62 ______________ now multiply the first equation by -2 (only the first equation -6x and 6x cancel out so now you are left with +18y=24 +20y=-62 ------------- now simplify

OpenStudy (anonymous):

just think of it as an addition problem like 

OpenStudy (anonymous):

do you get what I am saying ? @sherlocked1270

OpenStudy (anonymous):

well I know how to do a regular problem where the y values are the same but the y values are different this time

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

that's why you multiply the first equation by -2 to make the x values the same and once you found the y values you can substitute in and you both values